OpenStudy (anonymous):

Two urns contain white balls and yellow balls. The first urn contains 3 white balls and 6 yellow balls and the second urn contains 3 white balls and 8 yellow balls. A ball is drawn at random from each urn. What is the probability that both balls are white? A. B. C. D.

OpenStudy (anonymous):

A.2/33 B.1/9 C.1/11 D. 3/10

OpenStudy (cwrw238):

first event probability of choosing white ball = 3/9 = 1/3 second .......................................................... = 3/11 since the 2 events are independent the probabilities are multiplied so can you find the correct one?

OpenStudy (anonymous):

b.1/9

OpenStudy (anonymous):

Is that correct?

OpenStudy (cwrw238):

no whats 1 / 3 * 3/11

OpenStudy (cwrw238):

no whats 1 / 3 * 3/11

OpenStudy (cwrw238):



OpenStudy (cwrw238):

those are the 2 probabilities which are multiplied to get the required probability

OpenStudy (cwrw238):

those are the 2 probabilities which are multiplied to get the required probability

OpenStudy (cwrw238):

can you do that computation?

OpenStudy (cwrw238):

can you do that computation?

OpenStudy (anonymous):

I come out with 1/11

OpenStudy (cwrw238):

right - multiplying gives 3/33 which equals 1/11

OpenStudy (cwrw238):

right - multiplying gives 3/33 which equals 1/11

OpenStudy (anonymous):

thank you